VIBRONIC ANALYSIS FOR B̃ - X̃ TRANSITION OF ISOPROPOXY RADICAL

\underline{RABI CHHANTYAL-PUN} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus OH 43210.

Alkoxy radicals are important intermediates in combustion and atmospheric chemistry. Alkoxy radicals are also of significant spectroscopic interest for the study of Jahn Teller and pseudo Jahn Teller effects, involving the X̃ and Ã states. The Jahn Teller effect has been studied in methoxy. Substitution of one or two hydrogens by methyl groups transforms the interaction to a pseudo Jahn Teller effect in ethoxy and isopropoxy. Previously, moderate resolution scans have been obtained for B̃ - X̃ and B̃ - Ã transition systems, the latter observable at higher temperature. These measurements have shown that the X̃ and Ã states of isopropoxy are separated by only 60.7(7) cm^{−1} which indicates a strong pseudo Jahn Teller effect in the X̃ state. Such pseduo Jahn Teller coupling should also introduce additional bands into the B̃ - X̃ spectrum and a number of weaker transitions have been observed which may be caused by such effects. In this talk we present a vibronic analysis for the B̃ - X̃ transition based on the experimental results and also the results from recent quantum chemistry calculations.